******************************************************** ** ** ** IIII NNN NN GGGGG ** ** II NNNN NN GG GG LA PALMA ** ** II NN NN NN GG ** ** II NN NN NN GG GGG INFORMATION ** ** II NN NNNN GG GG ** ** IIII NN NNN GGGGG BULLETIN ** ** ** ** ** ** DATE: 1 September 1993 ** ** NUMBER: 0 ** ** ** ******************************************************** This is the first issue of the Isaac Newton Group - La Palma Information Bulletin; an e-mail news letter to keep the community of users of our telescopes informed on the actual situation on the mountain top. We invite readers to send comments (e-mail to BULLETIN@ING.IAC.ES), and to forward this bulletin to anyone who may be interested. Rene' Rutten ----------------------------------------------------------------------------- >>>>> WILLIAM HERSCHEL TELESCOPE <<<<< Chris Benn Downtime due to technical problems over the last 6 months has been 5%, about half the average over the last couple of years. A possible cause of the intermittent telescope oscillations over the last year may have been found; Martin Fisher discovered that one of the three transducers on the secondary mirror had failed. These record the tilt of the secondary as the telescope changes altitude, and send appropriate corrections to the altitude and azimuth drives. Several thermometers have been installed on the telescope tube and mirror. Temperatures are being measured and recorded throughout the day in order to improve our understanding of the thermal response of the telescope and its environment. The WYFFOS optical table was installed in the GHRIL enclosure in March. The fibre runs from prime focus will be installed later this summer. Prime-focus commissioning begins in late November. Work on creating an instrument test room (in the old optics lab) is in progress. This will provide facilities for the engineers to test each instrument off the telescope. Archiving is now possible to standard 10-inch tapes, exabytes or DATs, from either the instrument control computer LPVC or the data-reduction computer LPVE. >>>>> ISAAC NEWTON TELESCOPE <<<<< Emilios Harlaftis New turntable bearings for the Cassegrain rotator have recently been installed in order to improve the stability of the turntable rotation. This should also improve the telescope pointing. The dome suspension has been improved. As a result the oscillation of the telescope induced by the rotating dome is brought down to 2 arcseconds rms, compared to 5 arcseconds rms 3 years ago. However, the oscillation is still significant and additional improvements are under instigation. The data disks on the P-E instrument control computer are replaced and now provide a capacity which is four times larger. >>>>> INT PRIME FOCUS UNIT <<<<< Emilios Harlaftis Mechanical modifications to the Prime focus Unit have been made to reduce the stray light reaching the CCD. The Ford CCD system on the prime focus unit has been improved: the readout noise is now 12 electrons, the response is linear up to 60,000 counts, the horizontal charge-transfer is improved, and the underscan strip can now be read out. Discussions are underway to replace of the FORD chip with a thin one. >>>>> INTERMEDIATE DISPERSION SPECTROGRAPH <<<<< Emilios Harlaftis The FOS flat mirror has been replaced (end of May) and the IDS grating mechanism has been made more reliable. The flexure of IDS is less than 0.5 pixel/hour for zenith distances less than 60 degrees. >>>>> UNDERSAMPLING OF CCD IMAGES <<<<< Rene' Rutten / Vik Dhillon Undersampling of images (or spectra) on the CCD is a potential problem one should be aware of. When your image or spectrum has a width which is comparable to or smaller in size than a single pixel then subtle effects of CCD detectors are becoming important. In particular the fact that a single CCD pixel cannot be treated as a detector of uniform efficiency may affect your data. In a situation of severe undersampling the number of recorded counts will depend on the precise (sub-pixel) position of your stellar image or spectrum on the CCD. This will hamper accurate photometry or produce artificial ripples on your spectrum. Be aware of this problem, in particular when using FOS on the INT. For further information see the relevant articles in Gemini # 38 and 41. >>>>> TEK CCD <<<<< Rene' Rutten / Paul Whiteley The TEK1 CCD (1024 x 1024), which is currently in use on the WHT, has recently suffered from the following problems: (i) There have been intermittent shifts of one pixel in the spatial position of the images. These shifts occur seemingly at random. This was traced to a unreliable electronics module which has been replaced. (ii) The dark current has been too high, being about 60 ADU/hour. The most likely cause was poor cooling of the actual CCD in the cryostat. At the moment the dark current is down to ~5 electrons/hour. (iii) On-chip binning in the dispersion (X) direction results in non-linear response. This problem is still under investigation. For completeness we mention that the read-out noise on the TEK chip is ~5 ADU, and that the gain is 1.3 electrons/ADU. Furthermore TEK1 is known to produce a fringe pattern caused by its coating. This fringing starts to become important beyond ~6000 A. From 6500 A to 9000 A the fringe amplitude is about 2 to 3 %. From 9000 to 10000 A it becomes dramatically worse, going up to 10 and even 20 %. >>>>> CIRCULAR SPECTROPOLARIMETRY <<<<< Rene' Rutten The polarimetry module on the ISIS triple spectrograph behind the WHT is equipped not only with a halfwave plate for measuring linear polarisation, but also with a quarterwave plate for circular polarisation measurements. The equipment has recently been tested and is working fine. Since the quarterwave plate is tiny (taken from another instrument) the field of view is limited and only suitable for (near-) point sources. This also makes acquisition of objects somewhat cumbersome. For a full description of the polarisation module and its capabilities see `A users guide to WHT spectropolarimetry' by Tinbergen and Rutten. For further information contact Rene' Rutten or Vik Dhillon on La Palma. >>>>> SOFTWARE ON THE SUNS <<<<< Reynier Peletier Since the beginning of this year there is one Sun-Sparcstation in the control room of each telescope, and in the WHT terminal room. Observers can get guest accounts on these Sparc stations the same way as the guest accounts are set up on the VAX cluster. IRAF is running on the Sparcs and in the very near future we hope to also have Figaro and other Starlink packages available. Communication to the outside world through telnet is easy now. Furthermore, we are installing several programs, packages and utilities for observers to get the most out of their observing run. These include pgplot, xads, xgopher, tex and latex. For more information or suggestions, please contact Reynier Peletier at La Palma (peletier@ing.iac.es, ING::PELETIER). >>>>> WHO-IS-WHO FOR INSTRUMENT RESPONSIBILITIES ON SITE <<<<< Contact the following persons if you have queries about specific instruments (usernames given in brackets): WHT - Chris Benn (CRB) / Vik Dhillon (VSD) INT - Derek Jones (DHPJ) / Emilios Harlaftis (EHH) JKT - Phil Rudd (PJR) / Derek Jones (DHPJ) Imaging - Phil Rudd (PJR) / Reynier Peletier (PELETIER) INT prime - Emilios Harlaftis (EHH) / Reynier Peletier (PELETIER) ISIS/FOS2 - Rene' Rutten (RGMR) / Vik Dhillon (VSD) UES - Nic Walton (NAW) TAURUS - Nic Walton (NAW) LDSS - Mike Breare (BREARE) IDS/FOS1 - Emilios Harlaftis (EHH) / Derek Jones (DHPJ) RBS - Phil Rudd (PJR) PP - Derek Jones (DHPJ) / Phil Rudd (PJR) MPF - Rene' Rutten (RGMR) GHRIL - Vik Dhillon (VSD) / Mike Breare (BREARE) >>>>> ASTRONOMY PERSONNEL CHANGES <<<<< Peter Bunclark left the observatory on Aug. 1 Reynier Peletier started a tour as support astronomer on Aug. 1 Hans Slingerland will be leaving the observatory in late September. >>>>> SERVICE NIGHTS COMING UP <<<<< The following service nights are scheduled for the current semester Y: WHT Sep. 1 to 6 - half nights ISIS Sep. 20 - half night ISIS Oct. 29 - half night UES Nov. 8 - half night ISIS Nov. 15 - full night ISIS Dec. 18 - full night ISIS Dec. 20 - half night UES Jan. 5 - full night ISIS Nov. 14 - half night ISIS INT Sep. 7 - full night prime focus Sep. 15 - full night prime focus Oct. 12 - full night prime focus Nov. 22 - half night IDS Dec. 28 + 29 - full night prime focus Jan. 17 - full night IDS Jan. 24 - half night IDS JKT Sep. 12 - full night imaging Dec. 3 - full night imaging Dec. 11 + 12 - full night imaging Dec. 18 + 19 - full night imaging Jan. 1 + 2 - full night imaging Jan. 21 + 22 - full night imaging To apply for service time send you e-mail service application to SERVICE@AST-STAR.CAM.AC.UK >>>>> CONTACTING La Palma <<<<< La Palma can best be reached through INTERNET address: user@ING.IAC.ES In case of routing problems try: user@LPVE.ING.IAC.ES And through SPAN: 29146::user >>>>> NEXT patt APPLICATION DEADLINE <<<<< September 30, 1993 (for observations between Feb. 1 and Jul. 31, 1994). Note that applications are now required to have instrumental setup specified using the new Instrument Scheduling Form. ============================================================================== This information bulletin was prepared at the Isaac Newton Group of Telescopes. Distribution is by e-mail only. For free subscriptions send your Internet or JANET address to: >>>>>>>>>>>>> BULLETIN@ING.IAC.ES <<<<<<<<<<<<<<<< Or, for UK users, to: >>>>>>>>>>>>> DXC@CAST0.AST.CAM.AC.UK <<<<<<<<<<<<<<<< Suggestions regarding content are appreciated. ==============================================================================